OpenStudy (anonymous):

find the second partial derivative of z = y cos x So far I know that dz/dx is -sinx dz/dy, here im unsure, is it -y sin x??

OpenStudy (anonymous):

@OOOPS

OpenStudy (anonymous):

Nope, partial derivative w.r.t x, then y is constant, \[\dfrac{dz}{dx}=-ysinx\]

OpenStudy (anonymous):

\[\dfrac{dz}{dy}=cosx\]

OpenStudy (anonymous):

got me??

OpenStudy (anonymous):

yeah I think so, I will try for the second derative.

OpenStudy (anonymous):

you mean \(\dfrac{d^2z}{dx^2}\)?? second partial derivative??

OpenStudy (anonymous):

(dz/dx)^2 = -ycos x? (dz/dy)^2 = -sin x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\dfrac{d^2z}{dx^2}= -ycosx\] but \(\dfrac{d^2z}{dy^2}=0\) because to y, cos x is constant and (cosx)' =0

OpenStudy (anonymous):

oh okey, in my book it says -ycosx, -sin x and 0

OpenStudy (anonymous):

One more thing, pay attention at NOTATION, it is \(\dfrac{d^2z}{dx^2}\), NOT \((\dfrac{dz}{dx})^2\) they are different!!!

OpenStudy (anonymous):

You know why??

OpenStudy (anonymous):

oh i write the d^2z/dx^2 in my book but here I write like the wrong way, no I dont know why? please tell me:)

OpenStudy (anonymous):

\(\dfrac{dz}{dx}= -ysinx\rightarrow \dfrac{d^2z}{dx^2}= -ycosx\\while~~ (\dfrac{dz}{dx})^2 =(-ysinx)^2=y^2sin^2x\) See??

OpenStudy (anonymous):

oh, I didn't even realise that, thanx for making such equation. Btw where in the US did u study?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

So where di U study in USA?

OpenStudy (anonymous):

not did, do . I am currently student in the U.S

OpenStudy (anonymous):

oh which university?

OpenStudy (anonymous):

Temple university

OpenStudy (anonymous):

Hmm, never heard of it, where in US is it?

OpenStudy (anonymous):

http://temple.edu/

OpenStudy (anonymous):

cool it looks like a good university:)

OpenStudy (anonymous):

I don't know. I apply to that school because it is my neighbor.

OpenStudy (anonymous):

well atleast the education system is good there?

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

o'well, thanx for the time and help:)

OpenStudy (anonymous):

:)